OpenStudy (anonymous):

The difference between two square numbers is 19. What are the two square numbers?

OpenStudy (psymon):

There;s actually a pattern in the difference between two consecutive square numbers. So the first two squared numbers 1 and 4. 4 - 1 = 3 Next one 4 and 9. 9 - 4 = 5 9 and 16. 16 - 9 = 7 16 and 25. 25 - 16 = 9 25 and 36. 36 - 25 = 11 See a pattern? Bet you can then find the two that are 19 apart :3

OpenStudy (jdoe0001):

\(\begin{matrix} \text{Psymon}\\ \text{1 and 4. 4 - 1 =}&\text{3}\\ \text{Next one}\\ \text{4 and 9. 9 - 4 =}&\text{ 5}\\ \text{9 and 16. 16 - 9 =}&\text{ 7}\\ \text{16 and 25. 25 - 16 =}&\text{ 9}\\ \text{25 and 36. 36 - 25 =}&\text{ 11}\\ \end{matrix}\) See a pattern?

OpenStudy (anonymous):

OK, is there a formula such as 2n + 1 or anything like that?

OpenStudy (psymon):

I suppose you could put it like that, where n is the number you are squaring. That would give us as n =9, which of course the difference between 9^2 and 10^2 is the desired 19. That makes more sense like that :3

OpenStudy (anonymous):

Is the difference always a prime number?

OpenStudy (anonymous):

i am more confused now, i am 11 years old can you simplify your answer

OpenStudy (anonymous):

Or is the difference always the sum of the two numbers: 1+2 = 3 2+3 = 5 etc

OpenStudy (anonymous):

rionogal, the two numbers are 9 and 10 right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Let's see 9 x 9 = 81 and 10 x 10 = 100 100 - 81 = 19 right?

OpenStudy (anonymous):

ok i get it now, thanks

zepdrix (zepdrix):

oo neat pattern :o